# Aligning Text to Image in Diffusion Models is Easier Than You Think

### v3

@@ 4 Experiments
Implementation details. In our experiments, to confirm the flexibility of SoftREPA, we utilize various open-source text-to-image diffusion models, including Stable Diffusion 1.5, Stable Diffusion XL, and Stable Diffusion 3 to evaluate text-to-image alignment in both image generation and editing tasks. The batch size was set to 16, and we trained the model for fewer than 30,000 iterations using two A100 GPUs.
For Stable Diffusion 3, we set the length of the soft token to 4 and used 5 layers to attach the soft tokens for further experiments. In the case of Stable Diffusion 1.5 and Stable Diffusion XL, the soft token length was set to 8, and soft tokens were applied only to the Down and Middle block layers of the UNet’s original embedding. Refer to further implementation detail in Appendix B.
@@
Human Preference Text Alignment Image Quality
[…]
SD1.5 0.177 0.265 0.251 0.438
SD1.5 (Ours) 0.274 0.271 0.252 0.437
SDXL 0.750 0.267 0.273 0.420
SDXL (Ours) 0.852 0.268 0.283 0.423
SD3 0.942 0.263 0.280 0.424
SD3 (Ours) 1.085 0.269 0.289 0.428
Table 1: Quantitative results of image generation regarding image quality and text alignment of generated images.
We conducted text-to-image generation experiments on SD1.5, SDXL, and SD3, training soft tokens using the text-image paired COCO dataset [17]. The implementation details are provided in Appendix B. To evaluate the generated images, we assessed human preference scores [35, 12], text-image alignment [38, 34], and image quality [37, 27] on COCO-val 5K dataset. As shown in Tab. 1, the proposed method generally outperforms baseline approaches in both diffusion and rectified flow model regardless of the model architecture. In particular, detailed textual descriptions are more accurately reflected in the generated images, as demonstrated in the qualitative results (Fig. 3).
Figure 4: The qualitative results of text guided image editing comparing SD3 and SD3 with proposed method. The FlowEdit [14] is used as the editing method for both baseline and SoftREPA.
Figure 5: Quantitative comparisons between baseline editing and editing with soft tokens on DIV2K and Cat2Dog dataset. We evaluate CLIP score and LPIPS for each method. Lines connected with dots represent different CFG scales of the same method.
Human Preference Text Alignment Image Quality
[…]
[14] DIV2K 0.380 0.260 0.255 0.154
[14] (Ours) DIV2K 0.466 0.263 0.260 0.149
[14] Cat2Dog 0.937 0.225 0.266 0.199
[14] (Ours) Cat2Dog 1.111 0.226 0.269 0.173
Table 2: Quantitative evaluation of image editing performance in terms of image quality and target text alignment. The editing process follows FlowEdit [14], with the only difference being the use of soft tokens.
Figure 6: Additional qualitative results of text guided image editing on Cat2Dog dataset.
@@
To further demonstrate the improved text-image alignment achieved by our method, we conducted text-guided image editing experiments. These experiments are based on FlowEdit [14], an image editing framework for flow-matching models built on SD3. For the image editing dataset, we selected 800 high-quality images from the DIV2K dataset [1] and generated detailed captions for the source and target images using LLaVA [19]. The dataset generation process is described in detail in  Appendix C.
According to FlowEdit [14], the optimal CFG scale for the target text condition is set to 13.5, with 33 editing steps out of 50 total timesteps. In contrast, with our proposed soft tokens, the CFG scale is reduced to 11, and the number of editing steps is reduced to 30, suggesting that our method could provide a more efficient way to capture the joint text-image distribution. In Fig. 4, the edits in both single and multi-concept scenarios are more effectively reflected in the modified images compared to vanilla SD3. Additional qualitative results are conducted on Cat2Dog dataset Fig. 6 with target CFG scale as 7.
Furthermore, the incorporation of soft tokens leads to improvements in both text-image alignment and overall image quality (Fig. 5). The quantitative results of human preference scores, text alignment, and perceptual image quality evaluation show consistent enhancements with the use of soft tokens (Tab. 2). Detailed results for various parameter settings are provided in Appendix E.
Figure 7: Ablation study analyzing the effect of the number of layers incorporating soft text tokens and the length of tokens. The evaluation is conducted using ImageReward, CLIP score and LPIPS. Each metric is normalized within its respective range: ImageReward $[0.7,1.1]$, CLIP score $[0.25,0.28]$, and LPIPS $[0.42,0.44]$.
@@
We conducted an ablation study to assess the impact of both the number of layers incorporating soft text tokens and the length of these tokens on the text-to-image generation quality of SD3 using COCO val dataset [17] (Fig. 7). A value of 0 corresponds to the vanilla SD3 model, and the number of layers is counted from the first layer. Each comparison is conducted with either a fixed token length of 8 or a fixed number of 5 layers. Our findings indicate that incorporating soft tokens beyond layer $7$ severely degrades image generation quality. The optimal performance, aligned with human preferences (ImageReward [35]), was observed when soft tokens were applied within layers $2\sim 5$. This aligns with the model’s architectural design, where early layers primarily facilitate text-image alignment, while later layers focus on enhancing image fidelity [21]. However, text-image alignment scores (CLIP) continued to improve as more layers were tuned with soft tokens, which means that the model strongly follows the text conditions as more layers adopt the soft tokens while overall image quality remained stable across different layer configurations. The supporting qualitative results are presented in Appendix F.
Regarding token length, we found that using an excessive number of tokens, particularly beyond a length of 8, negatively impacted image generation quality. This suggests that longer token representations may lead to overfitting on the fine-tuning dataset (Fig. 7). Shorter token lengths ($1\sim 4$ tokens) were beneficial for preserving perceptual similarity to the source dataset (LPIPS), whereas moderate token lengths ($4\sim 8$ tokens) significantly enhanced text-image alignment, as indicated by CLIP scores and ImageReward. Qualitative results supporting these findings are presented in Appendix F.
@@ 5 Related Works
Several methods have been proposed to address the text-image misalignment. Training free approach such as CFG++ [6] addresses this problem as the off-manifold phenomenon of conventional CFG and reformulates it to a manifold-constrained CFG to get better text-to-image alignment. Other training-free approach, attend-and-excite [4], adjust attention activations to enhance text-image consistency. DOODL [32] applies guidance to end-to-end manner. Meanwhile, training-based approaches such as Diffusion-DPOK [33] introduce alignment losses during training to both enforce stronger text-conditioning and visual realism. Following Diffusion-DPO, denoising diffusion policy optimization(DDPO) [3] enhances the alignment using feedback from the pre-trained vision-language model. However, these methods require either extensive retraining or an additional human annotation dataset. In this paper, we use the prior knowledge of diffusion model itself to get better text and image alignment, proposing a novel SoftREPA on soft tokens.
@@ 6 Conclusion
Appendix A Information Theory on Diffusion Model
@@ Appendix B Implementation Details
The training implementation details are shown in  Tab. 3. For Stable diffusion 3, the length of soft token as $4$ and the number of layers as $5$ are used for the experiments. For Stable diffusion 1.5 and Stable diffusion XL, we used the length of soft token as $8$. Soft tokens are added to the original embedding only on the Down and Middle block layers of UNet. Since we found that the performance of the Stable Diffusion 1.5 degrades when soft tokens are initialized with a random distribution. Therefore, unconditional text embedding was utilized for initializing soft tokens.
In Stable diffusion 3, we optimize the soft tokens only using the contrastive score matching loss ($L_{\text{SoftREPA}}$). In Stable diffusion 1.5 and Stable diffusion XL, we optimize the soft tokens with the combination of the contrastive score matching loss function ($L_{\text{SoftREPA}}$) and the small weight of denoising score matching loss. The code is available at https://github.com/softrepa/SoftREPA.
Models lr wd batch size iterations token init optimizer lr scheduler
SD1.5 1e-3 1e-4 16 26,000 $\varnothing$ AdamW CosineAnnealingWarmRestarts
SDXL 1e-3 1e-4 16 30,000 $N(0,0.02)$ AdamW CosineAnnealingWarmRestarts
SD3 1e-3 1e-4 16 30,000 $N(0,0.02)$ AdamW CosineAnnealingWarmRestarts
@@ Appendix C Prompt for Editing dataset
”Describe the object and background in the image”
@@
”You are an AI assistant for generating paired text prompts for real image editing tasks. Your goal is to modify a given text description by replacing an object with other while strictly following these rules:
@@
  • 2) The replacement must be significantly different from the original concept but contextually appropriate. Avoid unrealistic substitutions (e.g., changing ”rabbit on grass” to ”rocket on grass”).
@@
  • 6) Do not change or add colors. Specifically, when modifying a building, change only the appearance, not the type of building (e.g., do not change ”building” to ”church” or ”lighthouse”).
  • 7) Modify only one feature at a time. If changing an object (e.g., ”starfish” to ”sea turtle”), do not alter its color, shape, or other attributes.
@@
Output: The image features a close-up of a brown fox with a blue nose. The fox is standing in a grassy field, and the background is blurred, creating a focus on the fox’s face. The fox’s ears are perked up, and its eyes are open, giving it a curious and attentive expression. The fox’s fur is brown, and the grass in the background is green, creating a natural and vibrant scene.”
Appendix D Additional Results on T2I Generation
The overall quantitative results are shown in Tab. 4. The additional qualitative results on SD1.5 and SDXL are shown in Fig. 8 and Fig. 9
Human Preference Text Alignment Image Quality
@@ Appendix D Additional Results on T2I Generation
SD1.5 30 7.0 0.177 21.478 0.264 0.250 24.598 0.438
SD1.5 +Ours 30 7.0 0.273 21.541 0.270 0.252 25.358 0.437
SDXL 30 7.0 0.750 22.389 0.267 0.273 24.697 0.420
SDXL +Ours 30 7.0 0.852 22.620 0.268 0.283 26.040 0.423
SD3 28 4.0 0.942 22.549 0.263 0.280 31.595 0.424
SD3 +Ours 28 4.0 1.085 22.550 0.269 0.289 36.214 0.428
Table 4: Quantitative results of text to image generation on SD1.5, SDXL, and SD3.
@@
Appendix E Additional Results on Image Editing
The overall quantitative results of DIV2K and Cat2Dog dataset on different target CFG scale are shown in Tab. 5 and Tab. 6, respectively. Additionally, qualitative results on different target CFG scale are shown in Tab. 5. The additional qualitative results across different number of editing steps are shown in Fig. 10
Figure 10: The additional qualitative results on editing using FlowEdit. The three different results adopting soft tokens (edited1-3) are vary on the number of editing steps, $27,29,30$ out of $50$ total timesteps, respectively.

### v4

@@ 4 Experiments
Implementation details. In our experiments, to confirm the flexibility of SoftREPA, we utilize various open-source T2I diffusion models, including Stable Diffusion 1.5, Stable Diffusion XL, and Stable Diffusion 3 to evaluate text and image alignment in both image generation and text-based image editing tasks. For the training of softs token, the batch size was set to 16 with less than 30,000 iterations using two A100 GPUs. For Stable Diffusion 3, we set the length of the soft token to 4 and used 5 layers to attach the soft tokens for further experiments. In the case of Stable Diffusion 1.5 and Stable Diffusion XL, soft tokens are applied only to the Down or Middle block layers of the UNet. Refer to further implementation details in appendix B.
@@
We conducted text-to-image generation experiments on SD1.5, SDXL, and SD3, training soft tokens using the text-image paired COCO dataset [22]. The implementation details are provided in appendix B. To evaluate the generated images, we assessed human preference scores [44, 17], text-image alignment [47, 43], and image quality [46, 33] on COCO-val 5K dataset. Furthermore, we conducted evaluation on GenEval [9], compared with RankDPO [16], which is fine-tuned via preference optimization. As shown in table 1, in COCO dataset, the proposed method generally outperforms baseline approaches in both diffusion and rectified flow models regardless of the model architecture. Additionally, in GenEval, SD3 with SoftREPA mostly outperforms RankDPO [16]. The drop in the counting metric can be primarily attributed to the tendency of diffusion models to interpret text alignment in a way that encourages the generation of multiple instances of the referenced object. Further analysis and potential mitigation strategies are discussed in appendix D. In fig. 3, detailed textual descriptions are more accurately reflected in the generated images, as demonstrated in the qualitative results.
Figure 4: The qualitative results of text guided image editing comparing on SD3 with the proposed method. The FlowEdit [19] is used as the editing method for both baseline and SoftREPA.
Figure 5: Quantitative comparison of baseline and SoftREPA on DIV2K and Cat2Dog editing using CLIP Score and LPIPS with various CFG scales.
Figure 6: Ablation study on the number of layers and token length for soft tokens. Metrics are normalized to: ImageReward $[0.7,1.1]$, CLIP score $[0.25,0.28]$, and LPIPS $[0.42,0.44]$.
Memory and Speed Comparison
In table 1, we report the peak GPU memory usage (PM) and per-image latency for models with and without soft tokens, including SD1.5, SDXL, and SD3. All models were evaluated in float16 precision. The use of soft tokens introduces no change in peak memory and results in only a negligible increase in latency.
@@
To further demonstrate the improved text-image alignment achieved by our method, we conducted text-guided image editing experiments using PnP [37], MasaCtrl [4] with DDIM [34], Direct Inversion [14], and FlowEdit [19]. FlowEdit employs SD3, while the other methods use SD1.5. For evaluation, we used PIEBench [14], DIV2K [1], and Cat2Dog datasets. For DIV2K editing dataset, we selected 800 high-quality images and generated detailed source and target captions using LLaVA [24]; dataset construction details are provided in appendix C. In table 2, the quantitative results of human preference, text alignment, and structure distance show consistent enhancements with the use of soft tokens without compromising structure and background fidelity.
Considering that FlowEdit performs text-based image editing, where performance heavily depends on text-image alignment, this result indicates improved efficiency in modeling the joint text-image distribution using the proposed method. Furthermore, our method with soft tokens achieves more efficient editing by reducing the number of editing steps to 30, compared to 33 in FlowEdit. As shown in fig. 6, incorporating soft tokens across various CFG scales consistently enhances text-image alignment and image quality. Figure 4 further demonstrates that performance has been improved on both single and multi-concept editing compared to vanilla SD3. Additional results and details are provided in appendix F.
@@
We conducted an ablation study on SD3 using the COCO val dataset [22] to evaluate the effects of the number of layers using soft text tokens and token length on generation quality (fig. 6). A value of 0 corresponds to vanilla SD3, and the number of layers is counted from the first. Comparisons were made using either a fixed token length of 8 or a fixed depth of 5 layers. Our findings indicate that incorporating soft tokens beyond layer $7$ severely degrades image generation quality. The optimal performance, aligned with human preferences (ImageReward [44]), was observed when soft tokens were applied within layers $2\sim 5$. This aligns with the architecture, where early layers handle text-image alignment and later ones enhance fidelity [26]. Notably, CLIP scores continued to improve as more layers used soft tokens, indicating stronger adherence to text prompts, while overall image quality remained stable. Supporting qualitative results are provided in appendix G. For token length, using more than 8 tokens negatively impacted quality, suggesting overfitting to the fine-tuning data. Shorter lengths (1–4 tokens) preserved perceptual similarity (LPIPS), while moderate lengths (4–8 tokens) significantly improved text-image alignment (CLIP and ImageReward). Qualitative examples are shown in appendix G.
@@ 5 Related Works
Several methods have been proposed to address the text-image misalignment. Training free approach such as CFG++ [7] addresses this problem as the off-manifold phenomenon of conventional CFG and reformulates it to a manifold-constrained CFG to get better text-to-image alignment. Other training-free approach, attend-and-excite [5], adjust attention activations to enhance text-image consistency. DOODL [41] applies guidance to end-to-end manner. Meanwhile, training-based approaches such as Diffusion-DPO [42] introduce alignment losses during training to both enforce stronger text-conditioning and visual realism. Following Diffusion-DPO, denoising diffusion policy optimization(DDPO) [3] enhances the alignment using feedback from the pre-trained vision-language model. RankDPO [16] presents fully synthetic, ranking-based preference datasets generated by reward models instead of humans. However, these methods require either extensive full-finetuning or an additional annotation dataset. In this paper, we use the prior knowledge of diffusion model itself to get better text and image alignment, proposing a novel SoftREPA on soft tokens.
To enable efficient prompt-level fine-tuning, VPT [13] introduced learnable tokens for vision transformers while keeping the backbone frozen. CoOp [49] and CoCoOp [48] proposed prompt learning methods for CLIP-like vision-language models by optimizing prompts to distinguish between a predefined set of classes in image classification tasks. In recent works, MinorityPrompt [38] and Optical-Prompt [27] learned sample-specific prompts to improve diversity and temporal coherence during sampling steps, respectively. In contrast to these approaches, which typically tailor prompts to individual samples or specific class sets, we propose learning general-purpose, sample-independent soft tokens. These tokens are optimized over a distribution of text-image pairs, $(\bm{X},\bm{Y})$, effectively guiding the propagation of text features to better align with the joint text-image distribution.
@@ 6 Conclusion
Limitation and Potential Negative Impacts. While soft tokens effectively enhance text-image alignment, they may risk overemphasizing textual guidance at the expense of faithfully preserving the user’s intent in the prompt. Future work could explore regularization techniques to mitigate this issue. Additionally, this work does not address potential risks such as data poisoning or adversarial attacks during soft token training. Although these concerns are beyond the scope of this paper, they merit consideration in future research. Moreover, our method inherits any potential negative impacts of the underlying backbone models.
@@
Appendix A Information Theoretical Analysis of Diffusion Models
@@ Appendix B Implementation Details
Architecture Details
For Stable Diffusion 3, the main experiments use soft tokens of length 4, applied across 5 layers. In Stable Diffusion 1.5, soft tokens of length 4 are applied only to the Down block layers of the UNet’s conditional text embeddings. For Stable Diffusion XL, soft tokens of length 8 are used, applied to both the Down and Middle block layers of the UNet’s conditional text embeddings, without incorporating timestep dependency.
Training Details
The hyperparameters used during training are listed in table 3. For Stable Diffusion 3, the soft tokens are optimized solely using the contrastive score matching loss ($L_{\text{SoftREPA}}$). In contrast, for Stable Diffusion 1.5 and Stable Diffusion XL, optimization combines the contrastive score matching loss with a small weighting of the denoising score matching loss. We observed that initializing the soft tokens with a random distribution led to performance degradation in Stable Diffusion 1.5. To address this, we initialized the soft tokens using unconditional text embeddings. The code is publicly available at https://github.com/softrepa/SoftREPA.
Models lr wd batch size (positive, negative) iterations token init optimizer lr scheduler
SD1.5 1e-3 1e-4 $32(4,28)$ 26,000 $\varnothing$ AdamW CosineAnnealingWarmRestarts
SDXL 1e-3 1e-4 $16(1,15)$ 30,000 $N(0,0.02)$ AdamW CosineAnnealingWarmRestarts
SD3 1e-3 1e-4 $16(4,12)$ 30,000 $N(0,0.02)$ AdamW CosineAnnealingWarmRestarts
@@
Inference Details
A detailed description of the image generation algorithm on MM-Dit is provided in algorithm 1. At each layer, a distinct set of soft tokens is prepended to the text features and used exclusively within that layer. These soft tokens are not carried over to subsequent layers; instead, new soft tokens are introduced or omitted as appropriate. Their primary role is to guide the text tokens toward better text-image alignment, particularly during the early stages of the model.
Memory Efficiency
For the memory efficiency metrics reported in table 1, values were measured by averaging results over 50 runs with an A100 GPU. Image resolution was set to $512\times 512$ for SD1.5 and $1024\times 1024$ for SDXL and SD3.
Algorithm 1 Image Generation with Soft Tokens in MM-DiT
1:Gaussian noise $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
2:Text $\bm{Y}\sim p_{\text{data}}$
3:Soft token $\mathbf{s}\sim\text{Embedding}(k,t)$
4:Number of layers $N$, Threshold layer $L$
5:Time steps $\{t_{T},t_{T-1},...,t_{0}\}$
6:Initialize $\bm{H}_{\text{img}}^{(0,T)}\leftarrow\mathbf{z}$
7:Initialize $\bm{H}_{\text{text}}^{(0,T)}\leftarrow\text{TextEncoder}(\bm{Y})$
8:$n=|\bm{H}_{\text{text}}^{(0,T)}|$
9:for $t$ in $\{t_{T},t_{T-1},...,t_{0}\}$ do
10:    for $l=1$ to $N$ do
11:         if $k\leq L$ then
12:             $\mathbf{s}^{(k,t)}\leftarrow\text{Embedding}(k,t)$
13:             $\hat{\bm{H}}_{\text{text}}^{(k-1,t)}\leftarrow[\mathbf{s}^{(k,t)};\bm{H}_{\text{text}}^{(k-1,t)}]$
14:         else
15:             $\hat{\bm{H}}_{\text{text}}^{(k-1,t)}\leftarrow\bm{H}_{\text{text}}^{(k-1,t)}$
16:         end if
17:         $\bm{H}_{\text{img}}^{(k,t)},\hat{\bm{H}}_{\text{text}}^{(k,t)}\leftarrow\text{Layer}_{l}(\bm{H}_{\text{img}}^{(k-1,t)},\hat{\bm{H}}_{\text{text}}^{(k-1,t)})$
18:         $\bm{H}_{\text{text}}^{(k,t)}\leftarrow\hat{\bm{H}}_{\text{text}}^{(k,t)}[-n:,:]$triangleright Drop soft tokens
19:    end for
20:end for
21:return $\mathbf{\hat{X}}=\text{Decoder}(\bm{H}_{\text{img}}^{(N,t_{0})})$
@@ Appendix C Prompt for Editing dataset
"Describe the object and background in the image"
@@
"You are an AI assistant for generating paired text prompts for real image editing tasks. Your goal is to modify a given text description by replacing an object with other while strictly following these rules:
@@
  • 2) The replacement must be significantly different from the original concept but contextually appropriate. Avoid unrealistic substitutions (e.g., changing "rabbit on grass" to "rocket on grass").
@@
  • 6) Do not change or add colors. Specifically, when modifying a building, change only the appearance, not the type of building (e.g., do not change "building" to "church" or "lighthouse").
  • 7) Modify only one feature at a time. If changing an object (e.g., "starfish" to "sea turtle"), do not alter its color, shape, or other attributes.
@@
Output: The image features a close-up of a brown fox with a blue nose. The fox is standing in a grassy field, and the background is blurred, creating a focus on the fox’s face. The fox’s ears are perked up, and its eyes are open, giving it a curious and attentive expression. The fox’s fur is brown, and the grass in the background is green, creating a natural and vibrant scene."
Appendix D Discussion on counting metric
Adopting Soft Tokens Only A Few Layers
As shown in section 4, our model demonstrates notable improvements in generating single and multiple objects, as well as in color and position-aware synthesis. However, it exhibits a relative decline in accurately generating the specified number of objects described in the text prompts. We hypothesize that this drop in the counting metric stems from the soft tokens excessively emphasizing textual cues, which can lead to overgeneration of object instances. To address this, we limited the use of soft tokens to the early layers of the model. As presented in table 4, applying soft tokens only to layers $1\sim 2$ preserves strong overall performance while mitigating the degradation in counting accuracy observed when soft tokens are applied across layers $1\sim 5$.
Incorporating Counting Loss during training
To further enhance counting fidelity, we trained SoftREPA with a lightweight object-counting loss. Specifically, we employ a mean squared error (MSE) loss between the predicted object count, which is derived from denoised images, and ground-truth labels using the lightweight object detection module, YOLOv8 [39]. This variant, shown in the last row of table 4, shows improved counting accuracy without sacrificing overall generation quality, suggesting that it effectively reduces the tendency to replicate objects due to textual overemphasis.
GenEval
@@ Appendix D Discussion on counting metric
SD3 0.68 0.99 0.86 0.56 0.85 0.27 0.55
SD3 (Ours) 1 0.70 0.99 0.91 0.51 0.89 0.31 0.58
SD3 (Ours) 1$\sim$2 0.69 1.00 0.89 0.50 0.88 0.34 0.56
SD3 (Ours) 1$\sim$5 0.70 1.00 0.95 0.29 0.92 0.34 0.68
SD3 (Ours) + count 1$\sim$5 0.69 0.99 0.88 0.59 0.86 0.25 0.55
Table 4: Additional quantitative comparison on the GenEval [9] benchmark. Bold indicates the best performance, and underline denotes the second best. "Layers" refers to the transformer layers where soft tokens are applied.
Appendix E Additional Results on T2I Generation
Additional qualitative results for SD1.5 and SDXL on the COCO-val5K dataset are presented in fig. 7 and fig. 8, respectively. Furthermore, qualitative examples on the GenEval benchmark are provided in fig. 9.
@@ Appendix E Additional Results on T2I Generation
Figure 9: The qualitative results of text-to-image generation comparing SD3 and SD3 with proposed method. The given text is from GenEval dataset.
Appendix F Additional Results on Image Editing
Qualitative comparisons of SoftREPA across various editing methods, including PnP [37], MasaCtrl [4] with DDIM [34], Direct Inversion [14], and FlowEdit [19], on the PIEBench dataset are presented in fig. 11, corresponding to the quantitative results in table 2.
Quantitative results on the DIV2K and Cat2Dog datasets under varying target CFG scales are reported in table 5 and table 6, respectively. The corresponding qualitative results for different CFG scales are also shown in table 5. Additionally, qualitative examples across different numbers of editing steps are provided in fig. 12 and fig. 10.